A leading process engineering company bought three (3) major pieces of equipment:


a shell and tube heat exchanger; a storage tank and a three-phase vertical separator.


The company purchased the exact same type of equipment three (3) times within a


ten (10) year period.


Due to audit concerns and missing receipts, you are tasked with determining the cost of EACH

piece of equipment in $US, using The Gauss-Jordan Matrix Method. You are only provided with the total cost(US$) of the following:

1st Time: 3 tanks + 4 heat exchangers + 2 vertical separators cost US$157,500


2nd Time: 2 tanks + 1 heat exchanger + 3 vertical separators cost US$90,000


3rd Time: 5 tanks + 3 heat exchangers + 2 vertical separators cost US$149,500




1
Expert's answer
2021-03-09T08:07:58-0500

Let a tank costs "x" $US, a heat exchanger costs "y" $US, and a vertical separator costs "z" $US. Then


"\begin{matrix}
 3x+4y+2z=157500 \\
 2x+y+3z=90000 \\
 5x+3y+2z=149500 \\

\end{matrix}"

Then


"A=\begin{pmatrix}
 3 & 4 & 2 & & 157500 \\
 2 & 1 & 3 & & 90000 \\
 5 & 3 & 2 & & 149500 \\
\end{pmatrix}"

Divide row 1 by 3 "(R_1=\dfrac{R_1}{3})"


"\begin{pmatrix}
 1 & 4/3 & 2/3 & & 52500 \\
 2 & 1 & 3 & & 90000 \\
 5 & 3 & 2 & & 149500 \\
\end{pmatrix}"

Subtract row 1 multiplied by 2 from row 2 "(R_2=R_2-(2)R_1)"


"\begin{pmatrix}
 1 & 4/3 & 2/3 & & 52500 \\
 0& -5/3 & 5/3 & & -15000 \\
 5 & 3 & 2 & & 149500 \\
\end{pmatrix}"

Subtract row 1 multiplied by 5 from row 3 "(R_3=R_3-(5)R_1)"


"\begin{pmatrix}
 1 & 4/3 & 2/3 & & 52500 \\
 0& -5/3 & 5/3 & & -15000 \\
 0 & -11/3& -4/3 & & -113000 \\
\end{pmatrix}"

Multiply row 2 by "-\dfrac{3}{5}" "(R_2=(-\dfrac{3}{5})R_2)"


"\begin{pmatrix}
 1 & 4/3 & 2/3 & & 52500 \\
 0 & 1 &-1 & & 9000 \\
 0 & -11/3& -4/3 & & -113000 \\
\end{pmatrix}"

Subtract row 2 multiplied by "\dfrac{4}{3}" from row 1 "(R_1=R_1-(\dfrac{4}{3})R_2)"


"\begin{pmatrix}
 1 & 0 & 2 & &40500 \\
 0 & 1 &-1 & & 9000 \\
 0 & -11/3& -4/3 & & -113000 \\
\end{pmatrix}"

Add row 2 multiplied by "\dfrac{11}{3}" to row 3 "(R_3=R_3+(\dfrac{11}{3})R_2)"


"\begin{pmatrix}
 1 & 0 & 2 & &40500 \\
 0& 1 &-1 & & 9000 \\
 0 & 0 & -5 & & -80000 \\
\end{pmatrix}"

Divide row 3 by "-5" "(R_3=\dfrac{R_3}{-5})"


"\begin{pmatrix}
 1 & 0 & 2 & &40500 \\
 0& 1 &-1 & & 9000 \\
 0 & 0 & 1 & & 16000 \\
\end{pmatrix}"

Subtract row 3 multiplied by 2 from row 1 "(R_1=R_1-(2)R_3)"


"\begin{pmatrix}
 1 & 0 & 0 & &8500 \\
 0& 1 &-1 & & 9000 \\
 0 & 0 & 1 & & 16000 \\
\end{pmatrix}"

Add row 3 to row 2 "(R_2=R_2+R_3)"


"\begin{pmatrix}
 1 & 0 & 0 & &8500 \\
 0& 1 & 0 & & 25000 \\
 0 & 0 & 1 & & 16000 \\
\end{pmatrix}"

A tank costs "8500" $US, a heat exchanger costs "25000" $US, and a vertical separator costs "16000" $US.